### v1

@@ 4 Numerical experiments and discussions
In this section, we provide numerical experiments for several linear inverse problems to complement the theoretical findings in Section 3. The experimental setting is identical to that in [13]. We employ three examples, i.e., s-phillips (mildly ill-posed), s-gravity (severely ill-posed) and s-shaw (severely ill-posed), which are generated from the code phillips, gravity and shaw, taken from the MATLAB package Regutools [7] (publicly available at http://people.compute.dtu.dk/pcha/Regutools/). All the examples are discretized into a finite-dimensional linear system with the forward operator $A_{\dagger}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ of size $n=m=1000$, with $A_{\dagger}x=(A_{{\dagger},1}x,\cdots,A_{{\dagger},n}x)$ for all $x\in\mathbb{R}^{m}$ and $A_{{\dagger},i}:\mathbb{R}^{m}\rightarrow\mathbb{R}$. To precisely control the regularity index $\nu$ in the source condition (cf. Assumption 2.1(ii)), we generate the exact solution $x_{\dagger}$ by
@@
with $x_{e}$ being the exact solution provided by the package and $\|\cdot\|_{\ell^{\infty}}$ the maximum norm of a vector. Note that the index $\nu$ in the source condition is slightly larger than the one used in (4.1) due to the existing regularity $\nu_{e}$ of $x_{e}$. The exact data $y_{\dagger}$ is given by $y_{\dagger}=A_{\dagger}x_{\dagger}$ and the noisy data $y^{\delta}$ is generated by $y^{\delta}_{i}:=y_{{\dagger},i}+\epsilon\|y_{\dagger}\|_{\ell^{\infty}}\xi_{i}$, $i=1,\cdots,n$, where $\xi_{i}$s follow the standard normal distribution, and $\epsilon>0$ is the relative noise level.
All the iterative methods are initialized to zero, with a constant step size $c_{0}=\|A_{\dagger}\|^{-2}$ for the Landweber method (LM) and $c_{0}=\mathcal{O}(c)$ for (r)SVRG, where $c=\min_{i}(\|A_{i}\|^{-2})=L^{-1}$. The constant step sizes $c_{0}$ is taken for rSVRG so as to achieve optimal convergence while maintaining computational efficiency across all noise levels. The methods are run for a maximum 1e5 epochs, where one epoch refers to one Landweber iteration or $nM/(n+M)$ (r)SVRG iterations, so that their overall computational complexity is comparable. The frequency $M$ of computing the full gradient is set to $M=2n$ as suggested in [15]. The operator $A$ for rSVRG is generated by the truncated SVD of $A_{\dagger}$ with $b=1/\big(1+2(\nu+\nu_{e})\big)$ and $a=(\|A_{\dagger}\|/\|y_{\dagger}\|)(n^{\nu+\nu_{e}}/c_{1})^{\frac{1}{1+2(\nu+\nu_{e})}}$, cf. Theorem 2.1 and Remark 3.1. Note the constant $c_{1}$ is fixed for each problem with different regularity indices $\nu$ and noise levels $\epsilon$. One can also use the randomized SVD to generate $A$.
@@
which can achieve order optimality. For rSVRG, $k_{*}$ is selected to be greater than the last index at which the iteration error exceeds that of LM upon its termination or the first index for which the iteration trajectory has plateaued. For SVRG, $k_{*}$ is taken such that the error is the smallest along the iteration trajectory. The accuracy of the reconstructions is measured by the relative error $e_{*}={\mathbb{E}[\|x_{k_{*}}^{\delta}-x_{\dagger}\|^{2}]^{\frac{1}{2}}}/{\|x_{\dagger}\|}$ for (r)SVRG, and $e_{*}=\|x_{k_{*}}^{\delta}-x_{\dagger}\|/\|x_{\dagger}\|$ for LM. The statistical quantities generated by (r)SVRG are computed based on ten independent runs.
@@
The numerical results for the examples with varying regularity indices $\nu$ and noise levels $\epsilon$ are presented in Tables 1, 2, and 3. It is observed that rSVRG achieves an accuracy (with much fewer iterations for relatively low-regularity problems) comparable to that for the LM across varying regularity. SVRG can also achieve comparable accuracy in low-regularity cases, indicating its optimality. However, with current step sizes, it is not optimal for highly regular solutions, for which smaller step sizes are required to achieve the optimal error [13]. Typically, problems with a higher noise level require fewer iterations. These observations agree with the theoretical results of Theorem 2.1 and Corollary 2.1. Moreover, the error of rSVRG at its plateau point is typically lower than that of the other two methods. The convergence trajectories of the methods for the examples with $\nu=0$ in Fig. 4.1 show the advantage of rSVRG over the other two methods as seen in Tables 1-3.
Table 1: The comparison between (r)SVRG and LM for s-phillips.
Method rSVRG ($c_{0}=c/4$) SVRG ($c_{0}=c/4$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 1.93e-2 102.825 1.17e-2 1.52e-2 1170.900 1.93e-2 758
5e-3 2.81e-2 14.325 2.52e-2 6.13e-2 137.625 2.81e-2 102
1e-2 3.79e-2 12.000 2.63e-2 7.93e-2 70.050 3.81e-2 68
5e-2 8.81e-2 6.075 4.58e-2 1.54e-1 11.100 9.44e-2 12
$0.25$ 1e-3 4.58e-3 206.700 4.29e-3 2.73e-2 819.225 4.58e-3 135
5e-3 1.48e-2 13.425 5.68e-3 5.73e-2 110.925 1.48e-2 60
1e-2 2.79e-2 12.825 9.43e-3 7.50e-2 58.650 2.81e-2 26
5e-2 4.13e-2 9.075 3.83e-2 1.37e-1 11.550 4.66e-2 10
$0.5$ 1e-3 2.87e-3 24.300 1.01e-3 2.73e-2 841.575 2.90e-3 94
5e-3 1.00e-2 12.675 3.79e-3 5.79e-2 115.050 1.21e-2 23
1e-2 1.33e-2 11.475 7.52e-3 7.53e-2 60.375 1.51e-2 16
5e-2 2.85e-2 9.150 2.49e-2 1.44e-1 12.675 2.92e-2 8
11 1e-3 1.53e-3 15.225 7.22e-4 2.76e-2 866.250 1.92e-3 25
5e-3 3.35e-3 17.775 3.28e-3 5.93e-2 163.800 3.44e-3 16
1e-2 5.36e-3 14.700 4.36e-3 7.76e-2 66.900 5.54e-3 12
5e-2 1.57e-2 12.075 1.57e-2 1.43e-1 11.850 1.82e-2 5
Table 2: The comparison between (r)SVRG and LM for s-gravity.
Method rSVRG ($c_{0}=c$) SVRG ($c_{0}=c$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 2.36e-2 279.525 1.30e-2 4.12e-2 1356.150 2.36e-2 1649
5e-3 3.99e-2 32.325 2.33e-2 9.05e-2 247.650 4.04e-2 255
1e-2 4.93e-2 25.425 3.65e-2 1.56e-1 93.900 5.30e-2 113
5e-2 8.56e-2 22.950 7.92e-2 3.50e-1 18.450 9.90e-2 22
$0.25$ 1e-3 6.16e-3 51.975 3.03e-3 4.74e-2 1550.400 6.50e-3 319
5e-3 1.56e-2 37.275 1.20e-2 1.25e-1 198.300 1.64e-2 71
1e-2 1.82e-2 27.150 1.27e-2 1.65e-1 164.325 2.32e-2 43
5e-2 5.12e-2 19.275 2.72e-2 4.05e-1 29.400 5.35e-2 12
$0.5$ 1e-3 3.34e-3 44.625 2.31e-3 3.82e-2 1106.400 3.39e-3 112
5e-3 7.56e-3 47.025 5.52e-3 1.26e-1 206.325 9.10e-3 40
1e-2 1.33e-2 44.550 1.04e-2 1.59e-1 176.100 1.41e-2 25
5e-2 3.38e-2 20.925 1.02e-2 4.00e-1 29.400 3.40e-2 8
11 1e-3 1.41e-3 48.000 9.87e-4 3.82e-2 1222.725 1.46e-3 42
5e-3 3.06e-3 35.400 1.11e-3 1.07e-1 259.800 4.11e-3 18
1e-2 3.17e-3 33.000 1.43e-3 1.57e-1 161.175 6.58e-3 12
5e-2 1.08e-2 23.175 8.15e-3 3.92e-1 29.400 1.48e-2 6
Table 3: The comparison between (r)SVRG and LM for s-shaw.
Method rSVRG ($c_{0}=c$) SVRG ($c_{0}=c$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 4.94e-2 39.825 4.94e-2 3.41e-2 4183.950 4.93e-2 22314
5e-3 9.22e-2 57.375 6.88e-2 4.93e-2 132.675 9.28e-2 4858
1e-2 1.53e-1 23.025 1.11e-1 5.98e-2 71.775 1.53e-1 642
5e-2 1.74e-1 20.925 1.71e-1 1.46e-1 26.925 1.78e-1 68
$0.25$ 1e-3 1.69e-2 90.450 1.09e-2 2.01e-2 745.500 1.69e-2 1218
5e-3 2.21e-2 36.000 2.20e-2 4.34e-2 79.800 2.24e-2 139
1e-2 2.46e-2 23.625 2.24e-2 6.99e-2 56.550 2.59e-2 99
5e-2 5.21e-2 15.000 3.20e-2 1.75e-1 20.775 7.02e-2 24
$0.5$ 1e-3 2.97e-3 42.075 2.84e-3 2.05e-2 598.725 3.16e-3 169
5e-3 7.80e-3 30.075 3.81e-3 5.17e-2 85.275 8.83e-3 78
1e-2 1.55e-2 21.075 5.89e-3 7.51e-2 56.175 1.69e-2 42
5e-2 4.63e-2 18.825 4.13e-2 1.97e-1 19.050 5.36e-2 16
11 1e-3 1.60e-3 40.875 5.63e-4 2.07e-2 225.300 1.80e-3 54
5e-3 5.16e-3 41.475 2.81e-3 5.60e-2 84.075 6.13e-3 25
1e-2 7.24e-3 28.650 6.31e-3 8.20e-2 55.125 1.18e-2 19
5e-2 4.79e-2 18.000 1.91e-2 2.12e-1 16.650 5.26e-2 6

### v2

@@ 4 Numerical experiments and discussions
In this section, we provide numerical experiments for several linear inverse problems to complement the theoretical findings in Section 2. The experimental setting is identical to that in [17]. We employ three examples, i.e., s-phillips (mildly ill-posed), s-gravity (severely ill-posed) and s-shaw (severely ill-posed), which are generated from the code phillips, gravity and shaw, taken from the MATLAB package Regutools [9] (publicly available at http://people.compute.dtu.dk/pcha/Regutools/). All the examples are discretized into a finite-dimensional linear system with the forward operator $A_{\dagger}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ of size $n=m=1000$, with $A_{\dagger}x=(A_{{\dagger},1}x,\cdots,A_{{\dagger},n}x)$ for all $x\in\mathbb{R}^{m}$ and $A_{{\dagger},i}:\mathbb{R}^{m}\rightarrow\mathbb{R}$. To precisely control the regularity index $\nu$ in the source condition in Assumption 2.1(ii), we generate the exact solution $x_{\dagger}$ by
@@
with $x_{e}$ being the exact solution provided by the package and $\|\cdot\|_{\ell^{\infty}}$ the maximum norm of a vector. Note that the smoothness index $\nu$ in the source condition is slightly larger than the one in (4.1) due to the existing regularity $\nu_{e}$ of $x_{e}$. The exact data $y_{\dagger}$ is given by $y_{\dagger}=A_{\dagger}x_{\dagger}$ and the noisy data $y^{\delta}$ is generated by
$y^{\delta}_{i}:=y_{{\dagger},i}+\epsilon\|y_{\dagger}\|_{\ell^{\infty}}\xi_{i},\quad i=1,\cdots,n,$
where the noise components $\xi_{i}$s follow the standard normal distribution, and $\epsilon>0$ is the relative noise level.
All the iterative methods are initialized to zero, with a constant step size $c_{0}=\|A_{\dagger}\|^{-2}$ for the Landweber method (LM) and $c_{0}=\mathcal{O}(c)$ for (r)SVRG and SGD, where $c=\min_{i}(\|A_{i}\|^{-2})=L^{-1}$. The constant step size $c_{0}$ is taken for rSVRG so as to achieve optimal convergence while maintaining computational efficiency across all noise levels. The methods are run for a maximum 1e5 epochs, where one epoch refers to one Landweber iteration, $nM/(n+M)$ (r)SVRG iterations, or $n$ SGD iterations, so that their overall computational complexity is comparable. The frequency $M$ of computing the full gradient is set to $M=2n$ as suggested in [20]. The operator $A$ for rSVRG is generated by the truncated SVD of $A_{\dagger}$ with $b=1/\big(1+2(\nu+\nu_{e})\big)$ and $a=(\|A_{\dagger}\|/\|y_{\dagger}\|)(n^{\nu+\nu_{e}}/c_{1})^{\frac{1}{1+2(\nu+\nu_{e})}}$, cf. Theorem 2.1 and Remark 3.1, in which $\nu_{e}$ is taken to be 0.25 for phillips, 0.1 for gravity, and 0 for shaw. Note the constant $c_{1}$ is fixed for each problem with different regularity indices $\nu$ and noise levels $\epsilon$. One can also use the randomized SVD to generate $A$. Note that the a priori choice of the parameters $a$ and $b$ is of theoretical interest only due to their dependence on the smoothness parameter $\nu+\nu_{e}$, which however is often unknown in practice.
@@
which can achieve order optimality. For rSVRG, $k_{*}$ is selected to be greater than the last index at which the iteration error exceeds that of LM upon its termination or the first index for which the iteration trajectory has plateaued. For SVRG and SGD, $k_{*}$ is taken such that the error is the smallest along the iteration trajectory. The accuracy of the reconstructions is measured by the relative error $e_{*}={\mathbb{E}[\|x_{k_{*}}^{\delta}-x_{\dagger}\|^{2}]^{\frac{1}{2}}}/{\|x_{\dagger}\|}$ for both (r)SVRG and SGD, and $e_{*}=\|x_{k_{*}}^{\delta}-x_{\dagger}\|/\|x_{\dagger}\|$ for LM. The statistical quantities generated by (r)SVRG are computed based on ten independent runs.
@@
The numerical results for the examples with varying regularity indices $\nu$ and noise levels $\epsilon$ are presented in Tables 1, 2, and 3. It is observed that rSVRG achieves an accuracy (with much fewer iterations for relatively low-regularity problems) comparable to that for the LM across varying regularity. SVRG can also achieve comparable accuracy in low-regularity cases, indicating its optimality. However, with the chosen step size schedule, it is not optimal for highly regular solutions, for which smaller step sizes are required to achieve the optimal error [17]. While SGD can achieve optimal accuracy with the fewest iterations for problems with low-regularity solutions, its accuracy is no longer optimal for high-regularity cases (e.g., $\nu=1$). Moreover, (r)SVRG exhibits reduced oscillations near the optimum compared to SGD when the noise level $\epsilon$ is relatively large, cf. Fig. 4.1. We refer interested readers to [17] for a more comprehensive comparison of the performance between SVRG and SGD under suitable step size schedules, as well as the numerical results that indicate the advantages of SVRG over LM. Typically, for all the problems and the four methods, the case with a higher noise level requires fewer iterations to reach the desired accuracy. These observations agree with the theoretical results of Theorem 2.1 and Corollary 2.1. Moreover, the error of rSVRG at its plateau point is typically lower than that of the other three methods. Indeed, the results by rSVRG can be much more accurate (by a factor of 10) than that by SVRG for many settings, as indicated by the results in the tables. Furthermore, the required number of iterations is smaller for rSVRG than that for SVRG. These results clearly show the potential computational advantage of rSVRG (provided that the truncation level $J$ in rSVRG can be chosen properly). The convergence trajectories of the methods for the examples with $\nu=0$ in Fig. 4.1 show the advantage of rSVRG over the other three benchmark methods as seen in Tables 1-3.
Table 1: The comparison between (r)SVRG, SGD and LM for s-phillips.
Method rSVRG ($c_{0}=c/4$) SVRG ($c_{0}=c/4$) SGD ($c_{0}=c/4$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 1.93e-2 102.825 1.17e-2 1.52e-2 1170.900 9.48e-3 79.200 1.93e-2 758
5e-3 2.81e-2 14.325 2.52e-2 6.13e-2 137.625 2.20e-2 13.400 2.81e-2 102
1e-2 3.79e-2 12.000 2.63e-2 7.93e-2 70.050 3.04e-2 3.350 3.81e-2 68
5e-2 8.81e-2 6.075 4.58e-2 1.54e-1 11.100 6.91e-2 1.950 9.44e-2 12
$0.25$ 1e-3 4.58e-3 206.700 4.29e-3 2.73e-2 819.225 6.22e-3 31.700 4.58e-3 135
5e-3 1.48e-2 13.425 5.68e-3 5.73e-2 110.925 1.26e-2 4.950 1.48e-2 60
1e-2 2.79e-2 12.825 9.43e-3 7.50e-2 58.650 1.78e-2 2.100 2.81e-2 26
5e-2 4.13e-2 9.075 3.83e-2 1.37e-1 11.550 6.47e-2 1.750 4.66e-2 10
$0.5$ 1e-3 2.87e-3 24.300 1.01e-3 2.73e-2 841.575 6.45e-3 32.250 2.90e-3 94
5e-3 1.00e-2 12.675 3.79e-3 5.79e-2 115.050 1.10e-2 3.250 1.21e-2 23
1e-2 1.33e-2 11.475 7.52e-3 7.53e-2 60.375 1.61e-2 2.100 1.51e-2 16
5e-2 2.85e-2 9.150 2.49e-2 1.44e-1 12.675 6.10e-2 0.600 2.92e-2 8
11 1e-3 1.53e-3 15.225 7.22e-4 2.76e-2 866.250 5.79e-3 33.400 1.92e-3 25
5e-3 3.35e-3 17.775 3.28e-3 5.93e-2 163.800 1.05e-2 3.250 3.44e-3 16
1e-2 5.36e-3 14.700 4.36e-3 7.76e-2 66.900 1.52e-2 2.100 5.54e-3 12
5e-2 1.57e-2 12.075 1.57e-2 1.43e-1 11.850 5.61e-2 0.350 1.82e-2 5
Table 2: The comparison between (r)SVRG, SGD and LM for s-gravity.
Method rSVRG ($c_{0}=c$) SVRG ($c_{0}=c$) SGD ($c_{0}=c$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 2.36e-2 279.525 1.30e-2 4.12e-2 1356.150 1.66e-2 51.150 2.36e-2 1649
5e-3 3.99e-2 32.325 2.33e-2 9.05e-2 247.650 3.49e-2 7.450 4.04e-2 255
1e-2 4.93e-2 25.425 3.65e-2 1.56e-1 93.900 4.67e-2 2.050 5.30e-2 113
5e-2 8.56e-2 22.950 7.92e-2 3.50e-1 18.450 1.13e-2 0.150 9.90e-2 22
$0.25$ 1e-3 6.16e-3 51.975 3.03e-3 4.74e-2 1550.400 7.52e-3 11.600 6.50e-3 319
5e-3 1.56e-2 37.275 1.20e-2 1.25e-1 198.300 1.89e-2 2.100 1.64e-2 71
1e-2 1.82e-2 27.150 1.27e-2 1.65e-1 164.325 3.21e-2 2.100 2.32e-2 43
5e-2 5.12e-2 19.275 2.72e-2 4.05e-1 29.400 1.04e-1 0.150 5.35e-2 12
$0.5$ 1e-3 3.34e-3 44.625 2.31e-3 3.82e-2 1106.400 6.14e-3 11.350 3.39e-3 112
5e-3 7.56e-3 47.025 5.52e-3 1.26e-1 206.325 1.77e-2 2.100 9.10e-3 40
1e-2 1.33e-2 44.550 1.04e-2 1.59e-1 176.100 3.05e-2 2.100 1.41e-2 25
5e-2 3.38e-2 20.925 1.02e-2 4.00e-1 29.400 1.20e-1 0.150 3.40e-2 8
11 1e-3 1.41e-3 48.000 9.87e-4 3.82e-2 1222.725 6.00e-3 11.350 1.46e-3 42
5e-3 3.06e-3 35.400 1.11e-3 1.07e-1 259.800 1.73e-2 2.100 4.11e-3 18
1e-2 3.17e-3 33.000 1.43e-3 1.57e-1 161.175 2.97e-2 2.100 6.58e-3 12
5e-2 1.08e-2 23.175 8.15e-3 3.92e-1 29.400 9.98e-2 0.150 1.48e-2 6
Table 3: The comparison between (r)SVRG, SGD and LM for s-shaw.
Method rSVRG ($c_{0}=c$) SVRG ($c_{0}=c$) SGD ($c_{0}=c$) LM
$\nu$ $\epsilon$ $e_{*}$ $k_{*}$ $\lim_{k\to\infty}e$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$ $e_{*}$ $k_{*}$
0 1e-3 4.94e-2 39.825 4.94e-2 3.41e-2 4183.950 3.41e-2 2732.000 4.93e-2 22314
5e-3 9.22e-2 57.375 6.88e-2 4.93e-2 132.675 4.70e-2 69.400 9.28e-2 4858
1e-2 1.53e-1 23.025 1.11e-1 5.98e-2 71.775 5.35e-2 45.400 1.53e-1 642
5e-2 1.74e-1 20.925 1.71e-1 1.46e-1 26.925 1.21e-1 10.400 1.78e-1 68
$0.25$ 1e-3 1.69e-2 90.450 1.09e-2 2.01e-2 745.500 4.27e-3 45.200 1.69e-2 1218
5e-3 2.21e-2 36.000 2.20e-2 4.34e-2 79.800 1.49e-2 11.500 2.24e-2 139
1e-2 2.46e-2 23.625 2.24e-2 6.99e-2 56.550 2.26e-2 7.300 2.59e-2 99
5e-2 5.21e-2 15.000 3.20e-2 1.75e-1 20.775 5.96e-2 0.900 7.02e-2 24
$0.5$ 1e-3 2.97e-3 42.075 2.84e-3 2.05e-2 598.725 4.73e-3 14.850 3.16e-3 169
5e-3 7.80e-3 30.075 3.81e-3 5.17e-2 85.275 1.26e-2 4.150 8.83e-3 78
1e-2 1.55e-2 21.075 5.89e-3 7.51e-2 56.175 1.72e-2 0.900 1.69e-2 42
5e-2 4.63e-2 18.825 4.13e-2 1.97e-1 19.050 6.19e-2 0.900 5.36e-2 16
11 1e-3 1.60e-3 40.875 5.63e-4 2.07e-2 225.300 5.31e-3 13.800 1.80e-3 54
5e-3 5.16e-3 41.475 2.81e-3 5.60e-2 84.075 1.28e-2 0.900 6.13e-3 25
1e-2 7.24e-3 28.650 6.31e-3 8.20e-2 55.125 1.75e-2 0.900 1.18e-2 19
5e-2 4.79e-2 18.000 1.91e-2 2.12e-1 16.650 6.68e-2 0.900 5.26e-2 6
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
phillips gravity shaw
In addition, to shed further insights into rSVRG, we have also implemented an alternative variant, denoted by $\mbox{rSVRG}_{*}$, which uses the SVD of $A$, as described in the introduction. The constant step size is taken to be $c_{0}=\mathcal{O}(c_{*})$ for $\mbox{rSVRG}_{*}$, with $c_{*}=\min_{1\leq i\leq J}(\sigma_{i}^{-2})$ ($J$ denotes the truncation level, and is much smaller than the data size $n$). The relevant numerical results are presented in Fig. 4.2, in which we compare the rSVRG in Algorithm 2 with rSVRG, where one epoch refers to $nM/(J+M)$ rSVRG iterations. Across all the four noise levels, we consistently observe that the iteration trajectories of the variant rSVRG exhibit notable oscillations during the initial phase, which can be rather pronounced. However, the oscillations nearly disappear when the iterates are getting closer to the convergence. Moreover, for relatively mildly ill-posed problems, rSVRG reaches the convergence slightly faster than rSVRG, but the final converged values are visually indistinguishable. Thus, the two variants (rSVRG and rSVRG) perform more or less comparably with each other.